# Dislocation correlations in GaN epitaxial films revealed by EBSD and XRD

### v1

@@ 3.1 X-ray diffraction
The thin black lines in Figs. 2(a,b) represent Monte Carlo simulations of the respective diffraction curves. As screw dislocations are very uncommon in GaN films grown by metal-organic chemical vapor deposition, we have assumed a density of $8\times 10^{8}$ cm-2 of mixed a+c-type dislocations for sample 1. The screening distance is taken to be $R=1$ µm. For sample 2, grown by molecular beam epitaxy, edge dislocations of a density of $5\times 10^{10}$ cm-2 with the screening distance $R=0.3$ µm are taken to simulate the asymmetric reflections. The symmetric Bragg reflections are simulated by screw dislocations with a density of $5\times 10^{8}$ cm-2, which is two orders of magnitude smaller than the density of edge dislocations. Their screening distance is taken as $R=0.7$ µm. The contribution of screw dislocations to the asymmetric XRD profiles of sample 2 is negligible and they are not considered in the further analysis.
@@
We fit both the measured and the Monte-Carlo-simulated XRD curves in the same way, following Ref. [30]. The symmetric reflections are fitted first to get the densities of screw dislocations, and their contributions are taken into account in the fits of the asymmetric reflections, which give the densities of the edge dislocations. Each reflection is fitted independently from the others. The two essential fitting parameters are the dislocation density $\varrho$ and the screening distance $R$. Figures 2(c,d) present by full symbols the values of $\varrho$ and $R$ obtained in the fits of the measured XRD curves, while open symbols are the results of the corresponding fits of the Monte-Carlo-simulated curves. The data are plotted as a function of the angle $\Psi$ between the diffraction vector and the surface. Thus, $\Psi=0$ corresponds to an in-plane reflection and $\Psi=90^{\circ}$ to a symmetric Bragg reflection. Plotting the fit parameters as a function of $\Psi$ is just a convenient way of presenting the data: ideally, no variation of the parameters is expected. The fits of the measured and simulated XRD profiles in different reflections from both samples show a small scatter in the values of the dislocation density $\varrho$ and a notably larger scatter in the values of $R$. This behavior results from the fact that the fit formula [30] accurately accounts for the orientation prefactor (a ‘contrast factor’) for the dislocation density. However, calculation of a respective prefactor for $R$ is considerably more difficult and has not been attempted here. Its absence does not affect the determination of the dislocation density, which was the primary aim of the fits, but introduces a scattering of the values of $R$. The EBSD study presented in the next section complements this reciprocal-space analysis by XRD with a real-space observation of the dislocation correlations, and allows for an accurate determination of the screening distance $R$. The values of $R$ obtained from EBSD maps are marked in Fig. 2(d) by dashed lines for comparison.
@@ 3.2 Electron backscatter diffraction: experiment
Figure 3: EBSD maps of the strain component $\varepsilon_{22}$ (a–c) in GaN samples 0, 1, and 2 and (d) in a Si(001) wafer, and (e,f) Monte Carlo simulation of the $\varepsilon_{22}$ maps for samples 1 and 2. Note that the strain obtained for Si(001) is an order of magnitude lower than in the GaN samples.
@@
Figures 3(a–c) present maps of the strain component $\varepsilon_{22}$ of the GaN samples 0, 1, and 2 obtained from the analysis of the EBSD measurements. Complete sets of maps of the six components of the strain tensor $\varepsilon_{ij}$ and the three components of the rotation tensor $\omega_{ij}$ are presented in the SM, Figs. S2, S3, and S4. As the EBSD maps are relative to an arbitrarily chosen reference value, we set the mean value of each map to zero. We follow the standard EBSD axes notation: the sample surface is the $xy$ plane, the $y$ axis is oriented such that the incident electron beam (inclined by $70^{\circ}$ with respect to the surface normal) is in the $yz$ plane. The sizes of the maps, $10\times 10$ µm2 for sample 1 and $2\times 2$ µm2 for sample 2, are chosen to cover a sufficiently large number of dislocations, of the order of $10^{3}$, for a statistical analysis.
@@
In a $2\times 2$ µm2 map of sample 0, just a single dislocation can be expected with a probability of 25%, compared to some $10^{3}$ dislocations in the maps of samples 1 and 2. So, sample 0 can be considered as dislocation free on the scale of the EBSD measurements. Nevertheless, the range of the strain variation in Fig. 3(a) is comparable with that in samples 1 and 2. In contrast to these latter samples, the strain map does not reveal correlations between strains in neighboring measurement points, which is quantified below by calculation of the strain–strain correlations. Since the strain range for sample 0 in Fig. 3(a) is an order of magnitude larger than the expected sensitivity of the EBSD measurements, the same measurement has been performed on a Si(001) wafer as presented in Fig. 3(d). The map also displays uncorrelated random noise, but the range of strain variations is of the order of $10^{-4}$, which is the expected sensitivity of high-resolution EBSD [13, 14] and an order of magnitude smaller than in all GaN samples under investigation. A whole set of the maps of all components of the strain and rotation tensor are presented in Fig. S5 in the SM and show the variations in the same range of $10^{-4}$.
@@
The map of $\varepsilon_{22}$ from sample 2 in Fig. 3(c) shows a notable difference between the $x$ and $y$ directions, with all features being extended in the $y$ direction. Figure S4 shows a similar anisotropy in the maps of all strain and rotation components of this sample. We have performed the same measurement with the sample rotated by 90 around its normal and found that the direction of the extension in the maps is also rotated by 90 to the new $y$ direction. A possible anisotropy of the sample is thus ruled out and it is established that the features in the maps are extended along the direction of the tilt of the incident electron beam. We show below that this extension is explained by a strongly anisotropic resolution of the EBSD measurements, and quantify this anisotropy.
@@
The shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ are of separate interest. Since the information depth of EBSD is small (less than 20 nm), the conditions of the stress-free surface $\sigma_{i3}=0\,(i=1,2,3)$, are satisfied. One of these conditions, $\sigma_{33}=0$, is used in the processing of the Kikuchi patterns, since only the differences of the strains $\varepsilon_{11}-\varepsilon_{33}$ and $\varepsilon_{22}-\varepsilon_{33}$ can be found from the positions of the Kikuchi bands and this condition allows the three normal strain components to be determined separately [16]. The other two conditions $\sigma_{i3}=0\,(i=1,2)$ are not used in the processing of the Kikuchi patterns. Since $\sigma_{i3}=2c_{44}\varepsilon_{i3}\,(i=1,2)$ in a hexagonal crystal, the strains $\varepsilon_{13}$ and $\varepsilon_{23}$ must be zero at the free surface. Figure 4(b) shows the probability distributions of $\varepsilon_{13}$ obtained from the respective EBSD maps of GaN samples 0, 1 and 2 as well as the silicon wafer. The probability distributions of the other shear strain component $\varepsilon_{23}$ are about two times narrower than those of $\varepsilon_{13}$. The maps of these components (shown in Figs. S2–S5 of the SM) exhibit uncorrelated random noise, as do all maps of the GaN sample 0 and the Si wafer. Figure 4(b) shows that all three GaN samples possess the same width of the $\varepsilon_{13}$ distribution, which is 16 times larger than the respective distribution for Si, and a factor of 2 narrower than the distribution of $\varepsilon_{22}$ from sample 0 in Fig. 4(a). Hence, the non-zero shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ are material specific. We postpone further discussion to Sec. 4.
@@
Figure 5: Strain–strain correlation functions of (a,b) samples 1 and 2 and (c,d) their Monte Carlo simulations.
@@
Figures 5(a,b) and 6(a,b) present the strain-strain correlations calculated from the maps of $\varepsilon_{22}$ in Figs. 3(a,b) and from similar maps of other strain components presented in Figs. S3 and S4. We calculate the correlation functions as averages $\left\langle\varepsilon(\mathbf{r}_{1})\varepsilon(\mathbf{r}_{2})\right\rangle$ over the respective maps. Figure 5 shows correlations in $x$ and $y$ directions (the difference $\mathbf{r}_{1}-\mathbf{r}_{2}$ is directed either along $x$ or $y$), while Fig. 6 shows correlations as a function of the distance $r=\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ averaged over all orientations of $\mathbf{r}_{1}-\mathbf{r}_{2}$. Since the strain of a single dislocation decays with the distance as $\varepsilon(\mathbf{r})\propto r^{-1}$, we expect the correlations to decay as $\ln\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ as long as $\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ is less than the distance $R$, at which the dislocation strain is screened by surrounding dislocations, and to become zero at larger distances. We therefore plot the correlations in Figs. 5 and 6 as a function of the logarithm of the distance. The expected linear decay of the correlations in this scale and a kink from the logarithmic decay to zero correlations are clearly seen in Figs. 6(a,b) and allow the determination of the screening distances: these distances are 2 µm for sample 1 and 0.3 µm for sample 2. This finding is the main result of the present paper.
@@
Figures 6(a,b) also show a drastic difference in the correlations between the in-plane strain components $\varepsilon_{11}$, $\varepsilon_{12}$, $\varepsilon_{22}$ and the shear strains $\varepsilon_{13}$, $\varepsilon_{23}$. The latter ones are expected to be zero due to zero stress conditions at the free surface but show distributions in Fig. 4(b) comparable in width to the in-plane strains. Their correlation functions in Figs. 6(a,b) are negligibly small compared to the correlations of the in-plane strains and quantify the visual impression of the respective maps in Figs. S3 and S4 as uncorrelated random noise.
@@ 3.3 Electron backscatter diffraction: Monte Carlo simulations
We simulate the maps of all strain and rotation components by the Monte Carlo method, as described in Sec. 2.4. We take the dislocation densities $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2 for samples 1 and 2, as determined in Sec. 2.1, and the mean distances between dislocations in the pairs of 2 µm and 0.3 µm, respectively, as derived above from the plots in Fig. 6(a,b). We simulate maps having the same areas and number of pixels as in the experiment, with the aim to obtain the same level of statistical errors due to averaging over the same number of dislocations. The strains are integrated over a 20 nm thick layer at the surface, corresponding to the information depth of EBSD measurements. The simulated maps of $\varepsilon_{22}$ in Fig. 3(e,f), as well as the simulated maps of all strain and rotation components shown in Figs. S6 and S7 show a good qualitative agreement with the measured maps.
@@
Figure 4(d) presents the probability distributions of the strain component $\varepsilon_{13}$ obtained from the Monte Carlo simulated maps of samples 1 and 2 in Figs. S6 and S7, respectively. This strain is equal to zero at the surface due to the strain-free boundary conditions, and the non-zero strains result from the integration of strains over the 20 nm-thick layer at the surface corresponding to the information depth of EBSD. The widths of the probability distributions of $\varepsilon_{13}$ in the simulated maps in Fig. 4(d) are at least by a factor of 5 smaller than the respective experimental widths in Fig. 4(b). Hence, the strains $\varepsilon_{13}$, $\varepsilon_{23}$ in the experimental maps are not the strains near the surface. They look, similarly to all maps of the dislocation-free sample 0, like uncorrelated random noise and may have the same origin, which is discussed in Sec. 4.4.
@@
The orientation-averaged correlation curves in Fig. 6(c,d), obtained from the Monte Carlo simulated maps, reveal the logarithmic decay of the strain correlations (straight lines in the coordinates of the figure) and a kink as a transition to zero correlations at distances exceeding the screening distance. The positions of the kinks coincide with the screening distances of 2 µm and 0.3 µm for samples 1 and 2 respectively, taken as an input of the Monte Carlo simulations. A gradual increase of correlations is seen at distances smaller than the resolution of 200 nm. Similar correlation functions for the different rotation components, both measured and Monte-Carlo simulated, are shown in Fig. S8 in the SM. They reveal the same logarithmic decay of correlations at distances smaller than $R$ and the absence of correlations at larger distances, where the characteristic kink is seen at the same distances $R$ as in the strain correlations of the respective samples.
@@
Figure 6: Orientation-averaged strain–strain correlation functions of (a,b) samples 1 and 2 and (c,d) their Monte Carlo simulations.

### v2

@@ 3.1 X-ray diffraction
The thin black lines in Figs. 2(a,b) represent Monte Carlo simulations of the respective diffraction curves. First, we simulate the symmetric Bragg reflections of sample 1 with screw dislocations. Agreement with the experimental diffraction profiles is obtained with a density of $8\times 10^{8}$ cm-2 and the screening distance of $R=1$ µm. Next, we turn to asymmetric reflections and include the edge components of the Burgers vectors in the calculations. Since screw dislocations are very uncommon in GaN(0001) films grown by metal-organic chemical vapor deposition [47, 48, 49, 50], we attribute the density of screw dislocations obtained in the symmetric reflections to the screw components of mixed a+c-type dislocations. The calculation of the diffraction profiles of asymmetric reflections with mixed dislocations agrees with the experimental diffraction curves, meaning that we do not need to add extra edge dislocations. We therefore conclude that sample 1 contains only mixed dislocations.
For sample 2 grown by molecular beam epitaxy, we also first simulate the symmetric Bragg reflections with screw dislocations. This revealed a screw dislocation density of $5\times 10^{8}$ cm-2 and a screening distance of $R=0.7$ µm. These dislocations provide only a small contribution to the widths of the asymmetric reflections. To simulate these reflections, we add edge dislocations with a density of $5\times 10^{10}$ cm-2, two orders of magnitude higher than the density of screw dislocations, with the screening distance of $R=0.3$ µm.
@@
We fit both the measured and the Monte-Carlo-simulated XRD curves in the same way, following Ref. [42]. The symmetric reflections are fitted first to get the densities of screw dislocations, and their contributions are taken into account in the fits of the asymmetric reflections, which give the densities of the edge dislocations. Each reflection is fitted independently from the others. The two essential fitting parameters are the dislocation density $\varrho$ and the screening distance $R$. Figures 2(c,d) present by full symbols the values of $\varrho$ and $R$ obtained in the fits of the measured XRD curves, while open symbols are the results of the corresponding fits of the Monte-Carlo-simulated curves. The data are plotted as a function of the angle $\Psi$ between the diffraction vector and the surface. Thus, $\Psi=0$ corresponds to an in-plane reflection and $\Psi=90^{\circ}$ to a symmetric Bragg reflection. Plotting the fit parameters as a function of $\Psi$ is just a convenient way of presenting the data: ideally, no variation of the parameters is expected.
The error bars representing the possible errors in determining the fit parameters from a fitted curve of a given reflection are smaller than the symbol size in Figs. 2(c,d) and can therefore be neglected. The scattering of the data points obtained from fitting the experimental curves in different reflections can have two sources, the accuracy of the fit function proposed in Ref. [42] and the presence of additional sources of strain, particularly a contribution from dislocations at the interface between the substrate and the relaxed epitaxial films. Our Monte Carlo simulations enable us to estimate the effect of the former by comparing the input parameters of the simulation with the output parameters of the fit.
For sample 1, the input parameters for the Monte Carlo simulation were the dislocation density $\varrho=8\times 10^{8}$ cm-2 of mixed dislocations and the average width of the dislocation pairs $R=1$ µm. Fitting the Monte Carlo simulated curves gives a dislocation density $\varrho=(8.3\pm 0.4)\times 10^{8}$ cm-2 and an average width of the dislocation pairs $R=1.4\pm 0.45$ µm. The errors represent the scattering between data points of different reflections in Figs. 2(c,d). For sample 2, the input parameters for the Monte Carlo simulations were the density $\varrho=5\times 10^{10}$ cm-2 of edge dislocations and $R=0.3$ µm. The fit of the simulated curves gives $\varrho=(5\pm 0.06)\times 10^{10}$ cm-2 and $R=(0.5\pm 0.05)$ µm. These results show that the fits following Ref. [42] quite accurately reproduce the dislocation density, but overestimate the screening distance by about 50%. This is not surprising since the fit formula for XRD curves contains prefactors for the dislocation density and for the screening distance. The prefactor for the dislocation density, also referred to as the “contrast factor” in powder diffraction, depends on the relative orientations of all the vectors involved in the problem (the diffraction vector, the Burgers vector, and the dislocation line direction). It can be accurately calculated in powder diffraction [53] and in diffraction from single crystals [42]. A respective prefactor for the screening distance has only been calculated analytically in the case of screw dislocations in the Wilkens’ model [5]. It can only be estimated for other cases, which limits the accuracy of determining the screening distance from the XRD profiles [12].
Fitting the experimental curves in different reflections yields for sample 1 a density $\varrho=(1\pm 0.3)\times 10^{8}$ cm-2 of mixed dislocations and $R=1\pm 0.7$ µm, and for sample 2 a density $\varrho=(5.8\pm 0.7)\times 10^{10}$ cm-2 of edge dislocations and $R=0.46\pm 0.14$ µm. These dislocation densities are larger than the densities of the spots in the cathodoluminescence or scanning electron microscopy images in Fig. 1 by factors of 2 for sample 1 and 3 for sample 2. We note that similar overestimates of the dislocation density in comparison with transmission electron microscopy results were obtained in previous studies [42, 54, 55].
The HR-EBSD study presented in the next section complements the reciprocal-space analysis by XRD with a real-space observation of the dislocation correlations. In this way, the screening distance $R$ can be determined more accurately and independent of the dislocation density determination. The values of $R$ obtained from the HR-EBSD maps are marked by dashed lines in Fig. 2(d) for comparison.
@@ 3.2 Electron backscatter diffraction: experiment
Figure 3: HR-EBSD maps of the strain component $\varepsilon_{22}$ (a–c) in GaN samples 0, 1, and 2 and (d) in a Si(001) wafer, and (e,f) Monte Carlo simulation of the $\varepsilon_{22}$ maps for samples 1 and 2. Note that the strain obtained for Si(001) is an order of magnitude lower than in the GaN samples. The $y$ axis is chosen in the direction of the tilt of the electron beam.
@@
Figures 3(a–c) present maps of the strain component $\varepsilon_{22}$ of the GaN samples 0, 1, and 2 obtained from the analysis of the HR-EBSD measurements. Complete sets of maps of the six components of the strain tensor $\varepsilon_{ij}$ and the three components of the rotation tensor $\omega_{ij}$ are presented in the SM, Figs. SM2, SM3, and SM4. As the HR-EBSD maps are relative to an arbitrarily chosen reference value, we set the mean value of each map to zero. We follow the standard HR-EBSD axes notation: the sample surface is the $xy$ plane, the $y$ axis is oriented such that the incident electron beam (inclined by $70^{\circ}$ with respect to the surface normal) is in the $yz$ plane. The sizes of the maps, $10\times 10$ µm2 for sample 1 and $2\times 2$ µm2 for sample 2, are chosen to cover a sufficiently large number of dislocations, of the order of $10^{3}$, for a statistical analysis.
@@
In a $2\times 2$ µm2 map of sample 0, just a single dislocation can be expected with a probability of 25%, compared to 500–700 dislocations in the maps of samples 1 and 2. So, sample 0 can be considered as dislocation free on the scale of the HR-EBSD measurements. Nevertheless, the range of the strain variation in Fig. 3(a) is comparable with that in samples 1 and 2. In contrast to these latter samples, the strain map does not reveal correlations between strains in neighboring measurement points, which is quantified below by calculation of the strain–strain correlations. Since the strain range for sample 0 in Fig. 3(a) is an order of magnitude larger than the expected sensitivity of the HR-EBSD measurements, the same measurement has been performed on a Si(001) wafer as presented in Fig. 3(d). The map also displays uncorrelated random noise, but the range of strain variations is of the order of $10^{-4}$, which is the expected sensitivity of HR-EBSD [25, 26] and an order of magnitude smaller than in all GaN samples under investigation. Whole sets of maps of all components of the strain and rotation tensors of Si(001) are presented in Fig. SM5 in the SM and show the variations in the same range of $10^{-4}$.
@@
The map of $\varepsilon_{22}$ from sample 2 in Fig. 3(c) shows a notable difference between the $x$ and $y$ directions, with all features being extended in the $y$ direction. Figure SM4 shows a similar anisotropy in the maps of all strain and rotation components of this sample. We have performed the same measurement with the sample rotated by 90 around its normal and found that the direction of the extension in the maps is also rotated by 90 to the new $y$ direction. A possible anisotropy of the sample is thus ruled out and it is established that the features in the maps are extended along the direction of the tilt of the incident electron beam. We show below that this extension is explained by a strongly anisotropic resolution of the HR-EBSD measurements, and quantify this anisotropy.
@@
The shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ are of separate interest. Since the information depth of EBSD is small (less than 20 nm), the conditions of the stress-free surface $\sigma_{i3}=0\,(i=1,2,3)$, are satisfied. One of these conditions, $\sigma_{33}=0$, is used in the processing of the Kikuchi patterns, since only the differences of the strains $\varepsilon_{11}-\varepsilon_{33}$ and $\varepsilon_{22}-\varepsilon_{33}$ can be found from the positions of the Kikuchi bands and this condition allows the three normal strain components to be determined separately [28]. The other two conditions $\sigma_{i3}=0\,(i=1,2)$ are not used in the processing of the Kikuchi patterns. Since $\sigma_{i3}=2c_{44}\varepsilon_{i3}\,(i=1,2)$ in a hexagonal crystal, the strains $\varepsilon_{13}$ and $\varepsilon_{23}$ must be zero at the free surface. Figure 4(b) shows the probability distributions of $\varepsilon_{13}$ obtained from the respective HR-EBSD maps of GaN samples 0, 1 and 2 as well as the silicon wafer. The probability distributions of the other shear strain component $\varepsilon_{23}$ are about two times narrower than those of $\varepsilon_{13}$. The maps of these components (shown in Figs. SM2–SM5 of the SM) exhibit uncorrelated random noise, as do all maps of the GaN sample 0 and the Si wafer. Figure 4(b) shows that all three GaN samples possess the same width of the $\varepsilon_{13}$ distribution, which is 16 times larger than the respective distribution for Si, and a factor of 2 narrower than the distribution of $\varepsilon_{22}$ from sample 0 in Fig. 4(a). Hence, the non-zero shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ are material specific. We postpone further discussion to Sec. 4.
@@
Figure 5: Strain–strain correlation functions of (a,b) samples 1 and 2 and (c,d) their Monte Carlo simulations. The vertical red arrows point out to the screening distances taken as input of the Monte Carlo simulations (2 µm and 0.3 µm for samples 1 and 2 respectively).
@@
Figures 5(a,b) and 6(a,b) present the strain-strain correlations calculated from the maps of $\varepsilon_{22}$ in Figs. 3(a,b) and from similar maps of other strain components presented in Figs. SM3 and SM4. We calculate the correlation functions as averages $\left\langle\varepsilon(\mathbf{r}_{1})\varepsilon(\mathbf{r}_{2})\right\rangle$ over the respective maps. Figure 5 shows correlations in $x$ and $y$ directions (the difference $\mathbf{r}_{1}-\mathbf{r}_{2}$ is directed either along $x$ or $y$), while Fig. 6 shows correlations as a function of the distance $r=\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ averaged over all orientations of $\mathbf{r}_{1}-\mathbf{r}_{2}$. Since the strain of a single dislocation decays with the distance as $\varepsilon(\mathbf{r})\propto r^{-1}$, we expect the correlations to decay as $\ln\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ as long as $\left|\mathbf{r}_{1}-\mathbf{r}_{2}\right|$ is less than the distance $R$, at which the dislocation strain is screened by surrounding dislocations, and to become zero at larger distances. We therefore plot the correlations in Figs. 5 and 6 as a function of the logarithm of the distance. The expected linear decay of the correlations in this scale and a kink from the logarithmic decay to zero correlations are clearly seen in Figs. 6(a,b) and allow the determination of the screening distances: these distances are 2 µm for sample 1 and 0.3 µm for sample 2. This finding is the main result of the present paper.
@@
Figures 6(a,b) also show a drastic difference in the correlations between the in-plane strain components $\varepsilon_{11}$, $\varepsilon_{12}$, $\varepsilon_{22}$ and the shear strains $\varepsilon_{13}$, $\varepsilon_{23}$. The latter ones are expected to be zero due to zero stress conditions at the free surface but show distributions in Fig. 4(b) comparable in width to the in-plane strains. Their correlation functions in Figs. 6(a,b) are negligibly small compared to the correlations of the in-plane strains and quantify the visual impression of the respective maps in Figs. SM3 and SM4 as uncorrelated random noise.
@@ 3.3 Electron backscatter diffraction: Monte Carlo simulations
We simulate the maps of all strain and rotation components by the Monte Carlo method, as described in Sec. 2.4. We take the dislocation densities $5\times 10^{8}$ cm-2 and $1.8\times 10^{10}$ cm-2 for samples 1 and 2, as determined in Sec. 2.1, and the mean distances between dislocations in the pairs of 2 µm and 0.3 µm, respectively, as derived above from the plots in Fig. 6(a,b). We simulate maps having the same areas and number of pixels as in the experiment, with the aim to obtain the same level of statistical errors due to averaging over the same number of dislocations. The strains are integrated over a 20 nm thick layer at the surface, corresponding to the information depth of EBSD measurements. The simulated maps of $\varepsilon_{22}$ in Fig. 3(e,f), as well as the simulated maps of all strain and rotation components shown in Figs. SM6 and SM7 show a good qualitative agreement with the measured maps.
@@
Figure 4(d) presents the probability distributions of the strain component $\varepsilon_{13}$ obtained from the Monte Carlo simulated maps of samples 1 and 2 in Figs. SM6 and SM7, respectively. This strain is equal to zero at the surface because of the strain-free boundary conditions. For edge dislocations, $\varepsilon_{13}$ and $\varepsilon_{23}$ are also equal to zero in the bulk of the crystal far from the surface. The non-zero strains in Figs. 4(d) and SM6, SM7 result from integrating the three-dimensional strain relaxation field near the surface over the 20 nm-thick surface layer corresponding to the EBSD information depth. The widths of the probability distributions of $\varepsilon_{13}$ in the simulated maps in Fig. 4(d) are at least 5 times smaller than the respective experimental widths in Fig. 4(b). Hence, the strains $\varepsilon_{13}$, $\varepsilon_{23}$ in the experimental maps are not the strains near the surface. They look, similarly to all maps of the dislocation-free sample 0, like uncorrelated random noise and may have the same origin, which is discussed in Sec. 4.4.
@@
The orientation-averaged correlation curves in Fig. 6(c,d), obtained from the Monte Carlo simulated maps, reveal the logarithmic decay of the strain correlations (straight lines in the coordinates of the figure) and a kink as a transition to zero correlations at distances exceeding the screening distance. The positions of the kinks coincide with the screening distances of 2 µm and 0.3 µm for samples 1 and 2 respectively, taken as an input of the Monte Carlo simulations. A gradual increase of correlations is seen at distances smaller than the resolution of 200 nm.
@@
Figure 6: Orientation-averaged strain–strain correlation functions of (a,b) samples 1 and 2 and (c,d) their Monte Carlo simulations. The vertical red arrows point out to the screening distances taken as input of the Monte Carlo simulations (2 µm and 0.3 µm for samples 1 and 2 respectively).
The probability distributions and the correlation functions for the different rotation components, both the measured and the Monte-Carlo simulated ones, are shown in Figs. SM8 and SM9 of the SM. Edge threading dislocations dominate in sample 2 and give rise to only in-plane rotations $\omega_{12}$ in the bulk of the sample. However, strain relaxation at the free surface results in out-of-plane rotations $\omega_{13}$ and $\omega_{23}$. These rotations are smaller than $\omega_{12}$ but larger than the shear strains $\varepsilon_{13}$ and $\varepsilon_{23}$ in Fig. 4(d), since the out-of-plane rotations $\omega_{13}$ and $\omega_{23}$ for edge dislocations are not zero at the surface. The correlation functions of rotations presented in Fig. SM9 show the same logarithmic decay of correlations at distances smaller than $R$ and the absence of correlations at larger distances. The characteristic kink is seen at the same distances $R$ as in the strain correlations of the respective samples.
The mixed dislocations used to model sample 1 give rise to equal widths of probability distributions of the in-plane and out-of-plane rotations in the Monte Carlo simulations in Fig. SM8(c). The experimental distribution of the in-plane rotation $\omega_{12}$ in Fig. SM8(a) is broader than the distributions of $\omega_{13}$ and $\omega_{23}$. In order to achieve agreement, edge dislocations with a density of $1\times 10^{9}$ cm-2 would need to be added. However, that disagrees with the dislocation density obtained from the cathodoluminescence image in Fig. 1 and also with the close widths of the XRD profiles in symmetric and asymmetric Bragg reflections in Fig. 2(a), which point out to mixed dislocations as a main part of the dislocation ensemble in sample 1. Therefore, only mixed dislocations are included in the Monte Carlo simulations of sample 1.
As we noted in Sec. 2.3, the measured EBSD maps suffer from sample drift, which is estimated to be below 18%. While the drift has no effect on the strain probability distributions, it can potentially affect the correlations of the strain tensor components, which are the main focus of our study. However, 18% drift refers to the relative displacement of the opposite corners of a measured map. Since we are looking for correlations between points separated by distances smaller than the screening distance $R$ (the correlations are absent at larger distances anyway), and since the map sizes are chosen to be an order magnitude larger than $R$, the effect of drift on separations smaller than $R$ is negligible. This is quantified in Fig. SM10 in the SM. In this example, we artificially expand a Monte Carlo-simulated map by 15% and compare the autocorrelation functions of the strain maps before and after the stretching. The difference between the curves can be neglected.
To estimate a possible statistical error in the screening distance determination, we collect in a single plot the autocorrelation functions obtained in measurements at different areas of the samples. They are compared with repeated Monte Carlo simulations performed using the same parameters, but with different values for the seed of the random number generator. These plots are compared in Fig. SM11 of the SM. Similar scattering of the curves is observed in both the experiment and the simulations. Therefore, the limited number of dislocations in a map area (only 500–700) rather than possible sample inhomogeneity is a main factor limiting the accuracy of the screening length determination. This could potentially be improved by measuring maps from sample areas containing larger number of dislocations. From Fig. SM11, the accuracy of our determination of the screening distance is estimated to be about 20%.